1.08(1.28)(8-x)=10
Answer by TimothyLamb(4379)

1.08(1.28)(8-x)=10
1.3824(8-x)=10
11.0592 - 1.3824x = 10
1.3824x = 11.0592 - 10
1.3824x = 1.0592
x = 1.0592 / 1.3824
x = 0.7662037
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php